# Rethinking Robustness in Machine Learning: A Posterior Agreement Approach

### v1

@@ Appendix C PA is not accuracy
It may be tempting to use Posterior Agreement to measure model performance instead of accuracy metrics. In Figure 10, we show an example of why this is a wrong use of PA. In particular, we display the performance of PA and of a classification model (DistilBERT-base) under three different shift strategies: (i) Levenshtein: addition, removal or substitution of characters in the sentence, (ii) Amplification: addition of adjectives reinforcing the sentence’s sentiment (iii) Contradiction: addition of adjective weakening the sentence’s sentiment. While in cases (i) and (iii) both metrics behave similarly, in case (ii) the addition of reinforcing adjectives has the effect of increasing the confidence in the model predictions, improving the overall performance. The sensibility of the method to this shift is correctly detected by PA, that penalizes the model decreasing its robustness score. This example again remarks that PA and accuracy are different metrics meant to investigate distinct aspects of a model.
@@ Appendix D Robustness and feature alignment
To better understand the impact of controlled domain shifts on robustness assessment through PA, we used ERM and IRM algorithms to train a ResNet18181818 model for $50$ epochs on $\mathcal{D}_{\mathrm{train}}$, using Adam with a learning rate of $10^{-2}$. From each validation sample $x^{\mathrm{val}}_{0},x^{\mathrm{val}}_{1}$, we selected $128$ observations to form a reduced validation set $\mathcal{D}_{\mathrm{sub}}=\{x^{\mathrm{sub}}_{0},x^{\mathrm{sub}}_{1}\}\subset\mathcal{D}_{\mathrm{val}}$. Both $x^{\mathrm{sub}}_{0}$ and $x^{\mathrm{sub}}_{1}$ share the same random seed $\tau^{\mathrm{val}}$, thus containing the same MNIST samples in the same order, and they differ only by a hue-based distribution shift (blue vs red; cf. Table 2).
@@ Appendix E Details on the DiagViB-6 Dataset
In this way, the training data is subject to both sampling randomness ($\tau_{0}^{\mathrm{train}}\neq\tau_{1}^{\mathrm{train}}$) and domain shift ($X_{0}\not\sim X_{1}$), while the validation and test splits each use a single instantiation, ensuring that distribution shift (i.e., changes in hue or other visual attributes) is the only significant source of randomness across $X^{\prime}$ and $X^{\prime\!\>\!\prime}$. Overall, we create two sets of 40 0004000040\,00040 000 images for training, two sets of 20 0002000020\,00020 000 images for validation, and six sets of 10 0001000010\,00010 000 images for testing.
@@
Table 2: Specific image factors associated with each environment considered in the model discriminability experiments. CC and LC account for centered center and centered low, respectively.
@@
First, when both $\mathcal{D}_{\mathrm{val}}^{(0)}$ and $\mathcal{D}_{\mathrm{val}}^{(1)}$ are drawn from source domains, two configurations are considered. If validation samples originate from the same distribution as training, they are in the same distribution (SD) setting. If they come from source domains but with different hue instantiations than the training data, they are in the in-distribution (ID) setting. Second, when $\mathcal{D}_{\mathrm{val}}^{(0)}$ is drawn from source domains and $\mathcal{D}_{\mathrm{val}}^{(1)}$ from target domains, the mixed distribution (MD) setting applies. Depending on the magnitude of the hue shift, we define two cases: a $1$-factor mixed distribution ($1$F-MD), where only hue varies, and a $5$-factor mixed distribution ($5$F-MD), where hue is combined with additional variations. Finally, when both $\mathcal{D}_{\mathrm{val}}^{(0)}$ and $\mathcal{D}_{\mathrm{val}}^{(1)}$ are drawn from target domains, the setup corresponds to the out-of-distribution (OOD) setting, where validation samples differ entirely from training.
This setup ensures that models are systematically tested under increasing levels of domain shift, from mild (ID, $1$F-MD) to severe ($5$F-MD, OOD). By doing so, we assess the robustness of model selection criteria when generalizing to unseen hue shifts. For more details on the final dataset configuration please refer to Table 3.
@@
Training 0 red dark CC large blank 1,4,7,9
1 blue dark CC large blank 1,4,7,9
Validation 0 red dark CC large blank 1,4,7,9
SD 1 red dark CC large blank 1,4,7,9
ID 1 blue dark CC large blank 1,4,7,9
1F-MD 1 magenta dark CC large blank 1,4,7,9
@@
Validation OOD 0 yellow dark CC large blank 1,4,7,9
1 magenta dark CC large blank 1,4,7,9
Table 3: Image factors associated with each of the environments considered in the model selection experiment. CC and UL account for ’centered center’ and ’upper left’, respectively.

### v2

@@ Appendix C PA is not accuracy
It may be tempting to use Posterior Agreement to measure model performance instead of accuracy measures. In Figure 10, we show an example of why this is a wrong use of PA. In particular, we display the performance of PA and of a classification model (DistilBERT-base) under three different shift strategies: (i) Levenshtein: addition, removal or substitution of characters in the sentence, (ii) Amplification: addition of adjectives reinforcing the sentence’s sentiment (iii) Contradiction: addition of adjective weakening the sentence’s sentiment. While in cases (i) and (iii) both measures behave similarly, in case (ii) the addition of reinforcing adjectives has the effect of increasing the confidence in the model predictions, improving the overall performance. The sensibility of the method to this shift is correctly detected by PA, that penalizes the model decreasing its robustness score. This example again remarks that PA and accuracy are different measures meant to investigate distinct aspects of a model.
@@ Appendix D Robustness and feature alignment
To better understand the impact of controlled domain shifts on robustness assessment through PA, we used ERM and IRM algorithms to train a ResNet1818 model for $50$ epochs on $\mathcal{D}_{\mathrm{train}}$, using Adam with a learning rate of $10^{-2}$. From each validation sample $x^{\mathrm{val}}_{0},x^{\mathrm{val}}_{1}$, we selected $128$ observations to form a reduced validation set $\mathcal{D}_{\mathrm{sub}}=\{x^{\mathrm{sub}}_{0},x^{\mathrm{sub}}_{1}\}\subset\mathcal{D}_{\mathrm{val}}$. Both $x^{\mathrm{sub}}_{0}$ and $x^{\mathrm{sub}}_{1}$ share the same random seed $\tau^{\mathrm{val}}$, thus containing the same MNIST samples in the same order, and they differ only by a hue-based distribution shift (blue vs red; cf. Table 2).
@@ Appendix E Details on the DiagViB-6 Dataset
In this way, the training data is subject to both sampling randomness ($\tau_{0}^{\mathrm{train}}\neq\tau_{1}^{\mathrm{train}}$) and domain shift ($X_{0}\not\sim X_{1}$), while the validation and test splits each use a single instantiation, ensuring that distribution shift (i.e., changes in hue or other visual attributes) is the only significant source of randomness across $X^{\prime}$ and $X^{\prime\!\>\!\prime}$. Overall, we create two sets of 40 00040\,000 images for training, two sets of 20 00020\,000 images for validation, and six sets of 10 00010\,000 images for testing.
@@
Table 2: Specific image factors associated with each environment considered in the model discriminability experiments. $C$ and LC account for centered center and centered low, respectively.
@@
First, when both $\mathcal{D}_{\mathrm{val}}^{(0)}$ and $\mathcal{D}_{\mathrm{val}}^{(1)}$ are drawn from source domains, two configurations are considered. If validation samples originate from the same distribution as training, they are in the same distribution (SD) setting. If they come from source domains but with different hue instantiations than the training data, they are in the in-distribution (ID) setting. Second, when $\mathcal{D}_{\mathrm{val}}^{(0)}$ is drawn from source domains and $\mathcal{D}_{\mathrm{val}}^{(1)}$ from target domains, the mixed distribution (MD) setting applies. Depending on the magnitude of the hue shift, we define two cases: a $1$-factor mixed distribution (11F-MD), where only hue varies, and a $5$-factor mixed distribution (55F-MD), where hue is combined with additional variations. Finally, when both $\mathcal{D}_{\mathrm{val}}^{(0)}$ and $\mathcal{D}_{\mathrm{val}}^{(1)}$ are drawn from target domains, the setup corresponds to the out-of-distribution (OOD) setting, where validation samples differ entirely from training.
This setup ensures that models are systematically tested under increasing levels of domain shift, from mild (ID, 11F-MD) to severe (55F-MD, OOD). By doing so, we assess the robustness of model selection criteria when generalizing to unseen hue shifts. For more details on the final dataset configuration, please refer to Table 3.
@@
Training 0 red dark $C$ large blank 1,4,7,9
1 blue dark $C$ large blank 1,4,7,9
Validation 0 red dark $C$ large blank 1,4,7,9
SD 1 red dark $C$ large blank 1,4,7,9
ID 1 blue dark $C$ large blank 1,4,7,9
1F-MD 1 magenta dark $C$ large blank 1,4,7,9
@@
Validation OOD 0 yellow dark $C$ large blank 1,4,7,9
1 magenta dark $C$ large blank 1,4,7,9
Table 3: Image factors associated with each of the environments considered in the model selection experiment. $C$ and UL account for ’centered center’ and ’upper left’, respectively.
Appendix F Autoattack results
In order to test our measure on more realistic benchmarks, we include an analysis of PA vs $\text{AFR}_{T}$ using the well-known AutoAttack library (Croce & Hein, 2020b).
F.1 CIFAR10
We test the previous models on the CIFAR-$10$ dataset under three attacks: APGD-CE & DLR (Croce & Hein, 2020b) and FAB Croce & Hein (2020a), all with $\ell_{\infty}=\ell/255$, $\ell\in[2,4,8,16,32]$. These attacks are significantly more effective than PGD and FMN, therefore, we are able to score robustness in the complete $0-100\%$ $\text{AFR}_{T}$ range.
Figure 14: PA (left) and $\text{AFR}_{T}$ (right) scores against increasing AR and $\ell_{\infty}$, for the APGD-CE attack of the Autoattack library.
Figure 15: PA (left) and $\text{AFR}_{T}$ (right) scores against increasing AR and $\ell_{\infty}$, for the APGD-DLR attack of the Autoattack library.
Figure 16: PA (left) and $\text{AFR}_{T}$ (right) scores against increasing AR and $\ell_{\infty}$, for the FAB attack of the Autoattack library.
In Figure 14-16, we can see the results for the attacks. The trends of the models are similar to the ones presented in the main paper, with $\text{AFR}_{T}$ failing to single out the undefended model. On the contrary, in the APGD family, PA scores its minimum already with $AR\ll 1$. Another notable phenomenon is represented by Addepalli et al. (2022) model, which is scored as the best or second-best model according to PA. In $\text{AFR}_{T}$, this happens only in the presence of large $AR$ and $\ell_{\infty}$ values, due to the faster worsening of the other models. This highlights once again the risk of inconsistency in the robustness evaluation when using $\text{AFR}_{T}$, while PA maintains an overall stable ranking across the $\ell_{\infty}$ values. As witnessed with PGD and FMN, PA correctly estimates the models’ performance and ranks them as $\text{AFR}_{T}$ ($AR=0$) does (this is not easily noticeable in the plots due to the scale used).
F.2 ImageNet
We select four of the fastest models and test them on the ImageNet dataset. Following (Croce & Hein, 2020b), we use the APGD-CE & APGD-DLR, an approximated, faster version of APGD-DLR for large datasets. We test with $\ell_{\infty}=\ell/255$, $\ell\in[2,4,8]$.
In Figure 17-18, we report the results. We obtain similar results to those in the previous case, with PA discriminating the weak models from the robust ones. In this case, JPEG+RS defense is only resisting the attack at higher power, and behaves as undefended for lower ones. This may be due to the different confidence in the outputs, on which the efficiency of the method depends. Otherwise, PA provides a robustness assessment that aligns with $\text{AFR}_{T}$ at high $AR$ values, showcasing its increased discriminability.
Figure 17: PA (left) and $\text{AFR}_{T}$ (right) scores against increasing AR and $\ell_{\infty}$, for the APGD-CE attack of the Autoattack library.
Figure 18: PA (left) and $\text{AFR}_{T}$ (right) scores against increasing AR and $\ell_{\infty}$, for the targeted APGD-DLR attack of the Autoattack library.